# A HHO formulation for variable density incompressible flows where the density is purely advected

### v1

@@ 3.5 H(div)-conformity and independence from irrotational body forces
in the definition of $sd^{{\rm slm}}_{T}$, see (8a) and (5). Note that the last two occurrences in the list above are required to make Neumann boundary conditions consistent with the reformulation and can be interpreted as adding and subtracting $\int_{\partial T^{\rm N}}\pi_{F}^{k+1}\psi\,\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$ to $sd^{{\rm slm}}_{T}$. Interestingly, the refactored HHO discretization, can be rewritten in terms of the modified pressure unknown $\widetilde{\underline{p}}_{h}\coloneqq\bigl(({p}_{T}-\pi_{T}^{k}\psi)_{T\in\mathcal{T}_{h}},({p}_{F}-\pi_{F}^{k+1}\psi)_{F\in\mathcal{F}_{h}}\bigr)$, being otherwise indistinguishable from the original version. This procedure allows to conclude that only the pressure field is affected by the irrotational part of body forces.
@@ 3.6 ESDIRK temporal discretization
$\bm{M}_{T}({\sf W}_{T})=\begin{bmatrix}\int_{T}{\rho}_{T}\;\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}&\int_{T}\bm{u}_{T}\;\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{z}}_{T}&\bm{0}\\ \bm{0}&\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T}&\bm{0}\\ \bm{0}&\bm{0}&\bm{I}\\ \end{bmatrix}$,
@@
where, $\bm{M}_{T}^{-1}({\sf W}_{T}){=}$
$\begin{bmatrix}\left(\int_{T}{\rho}_{T}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}\right)^{-1}&-\left(\int_{T}{\rho}_{T}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}\right)^{-1}\left(\int_{T}\bm{u}_{T}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{z}}_{T}\right)\left(\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T}\right)^{-1}&\bm{0}\\ \bm{0}&(\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T})^{-1}&\bm{0}\\ \bm{0}&\bm{0}&\bm{I}\\ \end{bmatrix}$.
@@ 3.7 Algebraic expression for Newton’s method and static condensation
${\sf\widehat{D}}_{i,T}^{st}(\underline{W}_{T}(t))$ $\coloneqq\frac{\delta_{i}{\sf W}_{T}^{!p}}{a_{ii}\,\delta t}+\bm{M}_{T}^{-1}({\sf W}_{T}|_{t+\delta t_{i}}){\sf D}_{T}^{s}(\underline{W}_{T}|_{t+\delta t_{i}})p$
$\qquad\qquad\qquad\qquad+\sum_{j=1}^{i-1}\frac{a_{ij}}{a_{ii}}\,\bm{M}_{T}^{-1}({\sf W}_{T}|_{t+\delta t_{j}}){\sf D}_{T}^{s}(\underline{W}_{T}|_{t+\delta t_{j}})$ (18a)
@@
$\sum_{T\in\mathcal{T}_{h}}\left({\sf\widehat{D}}^{st}_{i,\partial T}{-}\bm{J}_{\partial TT}\bigl(\bm{M}_{i,T}\bm{\widehat{J}}_{TT}\bigr)^{-1}\bigl(\bm{M}_{i,T}{\sf\widehat{D}}^{st}_{i,T}\bigr)\\ \right){=}\sum_{T\in\mathcal{T}_{h}}\bigl(\bm{S}_{T}\delta_{i}{\sf W}_{\partial T}\\ \bigr),$

### v3

@@ 3.5 H(div)-conformity and independence from irrotational body forces
Note that the last two occurrences in the list above are required to make Neumann boundary conditions consistent with the reformulation and can be interpreted as adding and subtracting $\int_{\partial T^{\rm N}}\pi_{F}^{k+1}\psi\,\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$ to $sd^{{\rm slm},\pi^{p}}_{T}$. Interestingly, the refactored HHO discretization, can be rewritten in terms of the modified pressure unknown $\widetilde{\underline{p}}_{h}\coloneqq\bigl(({p}_{T}-\pi_{T}^{k}\psi)_{T\in\mathcal{T}_{h}},({p}_{F}-\pi_{F}^{k+1}\psi)_{F\in\mathcal{F}_{h}}\bigr)$, being otherwise indistinguishable from the original version. This procedure allows to conclude that only the pressure field is affected by the irrotational part of body forces.
3.6 Flux formulation of the mass conservation equation
By means of the following integration by parts formula
$\int_{T}z_{T}\bm{u}_{T}\cdot\nabla\rho_{T}=\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{TF}\bigr){\rho}_{T}{z}_{T}-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla z_{T},$
it is straightforward to obtain the following equivalent reformulation of the steady mass conservation equation HHO formulation introduced in Sec. 3.4
$sd^{\rm sms}_{T,{\rm!bc}}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)=$ $-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla{z}_{T}+\frac{1}{2}\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)\,$
$+\int_{\partial T}\left(\frac{1}{2}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)({\rho}_{T}+\rho_{\partial T})+\frac{1}{2}\bigl|\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr|({\rho}_{T}-\rho_{\partial T})\right)({z}_{T}-{z}_{\partial T}).$ (18)
Notice that, in (18), $\int_{\partial T}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr){=}\int_{\partial T^{{\rm D},{\rm N}}}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)$, indeed $\rho_{\partial T}$ and ${z}_{\partial T}$ and are single valued over internal faces and $\bm{u}_{T}$ is H-div conforming. Introducing (18) in the definition of $sd^{\rm sms}_{T}$, see (11b), we finally get the following reformulation of the unsteady mass conservation HHO spatial discretization of (13): given $(\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T})\in\underline{\bm{V}}_{T}^{k}\times\underline{Z}_{T}^{k}\times\underline{Q}_{T}^{k+1}$, for all $\underline{z}_{T}\in\underline{Z}_{T}^{k}$,
$sd^{\rm mss}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)=$ $\int_{T}\frac{\partial\rho_{T}}{\partial t}z_{T}-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla{z}_{T}+\int_{\partial T}\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})({z}_{T}{-}{z}_{\partial T})$
$+\int_{\partial T^{\rm N}}\left(\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}+\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right){z}_{\partial T}$
$+\int_{\partial T^{\rm D}}\left(\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}+\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right){z}_{\partial T}.$
In the spatial discretization above, the numerical flux reads $\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\coloneqq\pi^{k}_{\partial T}\bigl(\phi^{\rm{td}}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})+\phi^{up}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\bigr)$, with
$\phi_{\partial T}^{\rm{td}}(\underline{\rho}_{T})$ $\coloneqq\widetilde{h}_{T}\left(\frac{\partial{\rho}_{T}}{\partial t}-\frac{\partial\rho_{\partial T}}{\partial t}\right),\;\rm{where}\;\widetilde{h}_{T}=\frac{h_{T}}{(k+1)^{2}};$
[…]
We remark that, since ${z}_{T}-{z}_{\partial T}\in\mathcal{P}^{k}(\partial T)$, we introduced the $L^{2}$-projection operator in the above definition of $\phi^{\rm mss}_{\partial T}$.
Let’s consider the requirement $r_{h}^{\rm mss}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h});\underline{z}_{h}\bigr)=0$ for each $(\underline{\bm{u}}_{h},\underline{\rho}_{h})\in\underline{\bm{V}}_{h}\times\underline{Z}_{h}^{k}$ and for all $\underline{z}_{h}\in\underline{Z}_{h}^{k}$ focusing on the following relevant choices of the test function
  1. 1.
    $\underline{z}_{h}=\left(({z}_{T})_{T\in\mathcal{T}_{h}},(0)_{F\in\mathcal{F}_{h}}\right)$, accordingly, $\forall T\in\mathcal{T}_{h}$ and $\forall{z}_{T}\in\mathbb{P}_{d}^{k}(T)$, we obtain the following local mass conservation statement
    $\int_{T}\frac{\partial\rho_{T}}{\partial t}z_{T}-\int_{T}\rho_{T}\bm{u}_{T}\cdot\nabla{z}_{T}+\int_{\partial T}\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T}){z}_{T}=0;$
  2. 2.
    $\underline{z}_{h}=\left((0)_{T\in\mathcal{T}_{h}},({z}_{F})_{F\in\mathcal{F}_{h}}\right)$, accordingly, we enforce the following continuity requirements for the numerical fluxes
    $\forall F\in\mathcal{F}_{h}^{\rm D},$ $\pi^{k}_{F}\left(\bigl(\bm{g}_{\rm D}{\cdot}\bm{n}_{TF}\bigr)^{\oplus}\rho_{\partial T}{+}\bigl(\bm{g}_{\rm D}{\cdot}\bm{n}_{TF}\bigr)^{\ominus}g_{\rm In}\right)=\pi^{k}_{F}\left(\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\right),$
    $\forall F\in\mathcal{F}_{h}^{\rm N},$ $\pi^{k}_{F}\left(\bigl(\bm{u}_{T}{\cdot}\bm{n}_{TF}\bigr)^{\oplus}\rho_{\partial T}{+}\bigl(\bm{u}_{T}{\cdot}\bm{n}_{TF}\bigr)^{\ominus}g_{\rm In}\right)=\pi^{k}_{F}\left(\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\right),$
    $\forall F\in\mathcal{F}_{h}^{\rm i},$ $\;\pi^{k}_{F}\left(\phi^{\rm mss}_{\partial T}(\bm{u}_{T}{,}\underline{\rho}_{T})\right)+\pi^{k}_{F}\left(\phi^{\rm mss}_{\partial T^{\prime}}(\bm{u}_{T^{\prime}}{,}\underline{\rho}_{T^{\prime}})\right)=0.$
    The above conditions can be rewritten to emphasize the upwind nature of the numerical flux
    $\forall F\in\mathcal{F}_{h}^{\rm D},$ $\pi^{k}_{F}\left(\widetilde{h}_{T}\frac{\partial{\rho}_{F}}{\partial t}{+}\left|\bm{u}_{T}{\cdot}\bm{n}_{TF}\right|\,{\rho}_{F}\right)=$ $\pi^{k}_{F}\left(\widetilde{h}_{T}\frac{\partial{\rho}_{T}}{\partial t}{+}\mathrm{upw}(\bm{g}_{\rm D},{\rho}_{T},g_{\rm In})\right),$
    $\forall F\in\mathcal{F}_{h}^{\rm N},$ $\pi^{k}_{F}\left(\widetilde{h}_{T}\frac{\partial{\rho}_{F}}{\partial t}{+}\left|\bm{u}_{T}{\cdot}\bm{n}_{TF}\right|\,{\rho}_{F}\right)=$ $\pi^{k}_{F}\left(\widetilde{h}_{T}\frac{\partial{\rho}_{T}}{\partial t}{+}\mathrm{upw}(\bm{u}_{T},{\rho}_{T},g_{\rm In})\right),$
    $\forall F\in\mathcal{F}_{h}^{\rm i},$ $\pi^{k}_{F}\left((\widetilde{h}_{T}{+}\widetilde{h}_{T^{\prime}})\frac{\partial{\rho}_{F}}{\partial t}{+}\left|\bm{u}_{T}{\cdot}\bm{n}_{TF}\right|{{\rho}_{F}}\right)=$ $\pi^{k}_{F}\left({\widetilde{h}_{T}}\frac{\partial{\rho}_{T}}{\partial t}{+}\widetilde{h}_{T^{\prime}}\frac{\partial{\rho}_{T^{\prime}}}{\partial t}{+}\mathrm{upw}(\bm{u}_{T},{\rho}_{T},{\rho}_{T^{\prime}})\right),$
    where
[…]
    |unTF|a, if   unTF0,
    |unTF|b, if   <unTF0.
[…]
Note that an internal face is such that $F=\partial T\cap\partial T^{\prime}$, with $^{\prime}\in\mathcal{T}_{h}, TTT\neq T^{\prime}$ and $\bm{n}_{TF}=-\bm{n}_{T^{\prime}F}$. We further remark that $\phi_{\partial T}^{\rm{up}}(\bm{u}_{T}{,}\underline{\rho}_{T})=0$ if the normal velocity component vanishes on a face $F\in\mathcal{F}_{T}$. In this event, thanks to the additional time derivative stabilization provided by $\phi_{\partial T}^{\rm{td}}(\underline{\rho}_{T})$, the face density is still well defined. A similar strategy was proposed to stabilize the linear momentum equations in the limit of vanishing viscosity, see [4].
3.7 Stability
In this section we tackle the stability of the HHO formulation in case of homogeneous Dirichlet boundary conditions and body force, i.e., $\partial\Omega_{{\rm D}}=\partial\Omega$, and $\bm{g}_{{\rm D}}=\bm{f}=0$. In order to show that, in the aforementioned setting, the HHO spatial discretizations are dissipative, we start by choosing $\underline{\bm{v}}_{T}=\underline{\bm{u}}_{T}$ in the steady linear momentum HHO discretization and $\underline{z}_{T}=\underline{\rho}_{T}$ in the steady mass HHO discretization, to get
$sd^{{\rm slm},\pi^{k}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{u}}_{T}\bigr)=\frac{1}{2}\int_{\partial T}\rho_{\partial T}\bigl|\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr|\left|\bm{\pi}_{T}^{k}\bm{u}_{T}-\bm{\pi}_{\partial T}^{k}\bm{u}_{\partial T}\right|^{2}$
$\qquad+\int_{T}\mu\,{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}:{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}+\int_{\partial T}\frac{\mu}{h_{F}}\,\left|\bm{\pi}_{\partial T}^{k}(\bm{u}_{T}-\bm{u}_{\partial T})\right|^{2}+\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}\underline{p}_{T}\cdot\bm{u}_{T}+\int_{\partial T^{\rm D}}\frac{\mu}{h_{F}}\,\left|\bm{u}_{\partial T}\right|^{2},$ (19)
$sd^{\rm sms}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{\rho}_{T}\bigr)=\frac{1}{2}\int_{\partial T}\bigl|\bm{u}_{T}\cdot\bm{n}_{TF}\bigr|\bigl({\rho}_{T}-\rho_{\partial T}\bigr)^{2}.$ (20)
We remark that, for the momentum equation spatial discretization, we consider $sd^{{\rm slm},\pi^{p}}_{T}$ with p=kp{=}k. In order to recast the HHO spatial discretizations as we did above, the following considerations are in order. The two terms appearing in lines (2)-(3), and the two terms in lines (7)-(8), are additive inverses if $\underline{\bm{v}}_{T}=\underline{\bm{u}}_{T}$ and if $\underline{z}_{T}=\underline{\rho}_{T}$, respectively. In case of homogeneous Dirichlet, according to (16), we get $\bm{u}_{T}\cdot\bm{n}_{\partial T}=\pi_{F}^{k+1}(\bm{g}_{D}{\cdot}\bm{n}_{TF})=0$. Accordingly, in (11b), all the terms involving integrals over Dirichlet boundaries can be disregarded. Following the same reasoning, most of the terms in (11a) can be omitted, notice in particular that the terms involving gradient reconstruction operators are skew-symmetric. As a consequence, the sole contribution arising from the weak imposition of Dirichlet BCs is the last term in (19).
The pressure-velocity coupling in (19) does not alter the energy balance, indeed we are able to infer
$\sum_{T\in\mathcal{T}_{h}}\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}\underline{p}_{T}\cdot\bm{u}_{T}=-\sum_{T\in\mathcal{T}_{h}}\int_{T}{p}_{T}\;\nabla\cdot\bm{u}_{T}+\sum_{T\in\mathcal{T}_{h}}\int_{\partial T}p_{\partial T}\bm{u}_{T}\cdot\bm{n}_{TF}=\sum_{T\in\mathcal{T}_{h}}\int_{\partial T^{\rm D}}p_{\partial T}\bm{u}_{T}\cdot\bm{n}_{TF}=0,$ (21)
where, in the first step, we introduced the definition of the gradient reconstruction operator, and in the second step, we used the fact that the element velocity $\bm{u}_{T}$ is divergence-free and H-div conforming, see Sec. 3.5, together with the single-valuedness of $p_{\partial T}$ on internal faces. The non-dissipation property follows, since, in case of homogeneous Dirichlet BCs, $\bm{u}_{T}\cdot\bm{n}_{TF}=\pi_{F}^{k+1}(\bm{g}_{D}{\cdot}\bm{n}_{TF})=0$. Accordingly, we remark that all the terms on the right hand side of (19)-(20) but $\int_{T}\mu\,{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}:{{\mathcal{G}}_{T}^{k}}\underline{\bm{u}}_{T}$ are non-negative. As demonstrated in [13, Ch. 5.1.6], see also [11, Remark 2.2], the HHO discretization of the Laplace operator introduced in (5) is coercive i.e., for all $T\in\mathcal{T}_{h}$ and all $\underline{\bm{v}}_{T}\in\underline{\bm{V}}_{T}^{k}$, it holds
$\int_{T}\,{{\mathcal{G}}_{T}^{k}}\underline{\bm{v}}_{T}:{{\mathcal{G}}_{T}^{k}}\underline{\bm{v}}_{T}+\int_{\partial T}\frac{1}{h_{F}}\,\left|\bm{\pi}_{\partial T}^{k}(\bm{v}_{T}-\bm{v}_{\partial T})\right|^{2}\gtrsim\int_{T}\,\nabla\bm{v}_{T}:\nabla\bm{v}_{T}+\int_{\partial T}\frac{1}{h_{F}}\,\left|(\bm{v}_{T}-\bm{v}_{\partial T})\right|^{2},$
where the hidden (positive) constant is independent of $h$ and $T$. As a consequence, we are able to conclude that both $sd^{{\rm slm},\pi^{k}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{u}}_{T}\bigr)$ and $sd^{\rm sms}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{\rho}_{T}\bigr)$ are non-negative.
Let’s now consider the unsteady linear momentum HHO discretization and the unsteady mass HHO discretization defined in (12) and (13), respectively. In order to reformulate the time derivative in the momentum equation, we introduce the following identity
$\left({\rho}_{T}\frac{\partial\bm{u}_{T}}{\partial t}+\frac{\bm{u}_{T}}{2}\frac{\partial{\rho}_{T}}{\partial t}\right)\cdot\bm{u}_{T}=\frac{1}{2}\frac{\partial(\rho\bm{u}_{T}\cdot\bm{u}_{T})}{\partial t}=\frac{1}{2}\frac{\partial\left|\sqrt{\rho}\bm{u}_{T}\right|^{2}}{\partial t}.$ (22)
Setting $\underline{\bm{v}}_{T}=\underline{\bm{u}}_{T}$ in $sd^{{\rm mnt},\pi^{k}}_{T}$ and $\underline{z}_{T}=\underline{\rho}_{T}$ in $sd^{\rm mss}_{T}$, using (22) and summing over all mesh elements, we arrive at the following stability argument
$\frac{d}{dt}\sum_{T\in\mathcal{T}_{h}}\int_{T}\bigl|\sqrt{\rho_{T}}\bm{u}_{T}\bigr|^{2}$ $=-2\sum_{T\in\mathcal{T}_{h}}sd^{{\rm slm},\pi^{k}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{u}}_{T}\bigr)\leq 0,$
$\frac{d}{dt}\sum_{T\in\mathcal{T}_{h}}\left(\int_{T}\bigl(\rho_{T}\bigr)^{2}+\frac{h_{T}}{(k+1)^{2}}\int_{\partial T}\bigl({\rho}_{T}-\rho_{\partial T}\bigr)^{2}\right)$ $=-2\sum_{T\in\mathcal{T}_{h}}sd^{\rm sms}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{\rho}_{T}\bigr)\leq 0.$ (23)
@@ 3.8 ESDIRK temporal discretization
$\bm{M}_{T}({\sf W}_{T})=\begin{bmatrix}\int_{T}{\rho}_{T}\;\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}&\int_{T}\frac{\bm{u}_{T}}{2}\;\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{z}}_{T}&\bm{0}\\ \bm{0}&\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T}&\bm{0}\\ \bm{0}&\bm{0}&\bm{I}\\ \end{bmatrix}$,
@@
where
$\bm{M}_{T}^{-1}({\sf W}_{T})=\begin{bmatrix}\left(\int_{T}{\rho}_{T}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}\right)^{-1}&-\left(\int_{T}{\rho}_{T}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{v}}_{T}\right)^{-1}\left(\int_{T}\frac{\bm{u}_{T}}{2}\widetilde{\sf{v}}_{T}\otimes\widetilde{\sf{z}}_{T}\right)\left(\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T}\right)^{-1}&\bm{0}\\ \bm{0}&(\int_{T}\widetilde{\sf{z}}_{T}\otimes\widetilde{\sf{z}}_{T})^{-1}&\bm{0}\\ \bm{0}&\bm{0}&\bm{I}\\ \end{bmatrix}$.
@@ 3.9 Algebraic expression for Newton’s method and static condensation
${\sf\widehat{D}}_{i,T}^{st}(\underline{W}_{T}(t))$ $\coloneqq\frac{\delta_{i}{\sf W}_{T}^{!p}}{a_{ii}\,\delta t}+\bm{M}_{T}^{-1}({\sf W}_{T}|_{t+\delta t_{i}}){\sf D}_{T}^{s}(\underline{W}_{T}|_{t+\delta t_{i}})+\sum_{j=1}^{i-1}\frac{a_{ij}}{a_{ii}}\,\bm{M}_{T}^{-1}({\sf W}_{T}|_{t+\delta t_{j}}){\sf D}_{T}^{s}(\underline{W}_{T}|_{t+\delta t_{j}})$ (29a)
@@
$\sum_{T\in\mathcal{T}_{h}}\left({\sf\widehat{D}}^{st}_{i,\partial T}{-}\bm{J}_{\partial TT}\bigl(\bm{M}_{i,T}\bm{\widehat{J}}_{TT}\bigr)^{-1}\bigl(\bm{M}_{i,T}{\sf\widehat{D}}^{st}_{i,T}\bigr)\right){=}\sum_{T\in\mathcal{T}_{h}}\bigl(\bm{S}_{T}\delta_{i}{\sf W}_{\partial T}\bigr),$